Meena kumari goes to a shop and buys a saree, costing ₹ 5225, including sales tax of 12%. The shopkeeper gives her discount, so that the price is decreased by an amount equivalent to sales tax. The price is decreased by ?
Answer by mananth(16946)

Meena kumari goes to a shop and buys a saree, costing ₹ 5225, including sales tax of 12%. The shopkeeper gives her discount, so that the price is decreased by an amount equivalent to sales tax. The price is decreased by ?

If Rs. 100 is the cost price the sales price will be Rs. 112
IfRs, 5225 is the sales price then cost price will be
(100/112) * 5225 = 4665.17
Discount = sales tax amount = 5225- 4665.17
5225-4665.17 = 559.83 which is the discount